# MDBench: Benchmarking Data-Driven Methods for Model Discovery

### v1

@@ 4.2 Performance on PDE Data
Compared to ODEs, PDE systems present a more challenging testbed due to the presence of spatial derivatives and higher-order terms, which tend to amplify noise. Figure 2(b) shows the aggregate NMSE and complexity of representative methods over a selected subset of PDE datasets for which all methods completed successfully (refer to Figure 10(b) for all methods). The selected datasets include Advection, Burgers, KdV, KS, Heat (Solar) 1D, and AD. The primary reasons for a method’s failure to discover an underlying model include: a lack of support for higher-dimensional data in the published software, failing to finish training before the timeout threshold was reached, and runtime errors within the published codebases. EQL and End2end are excluded from the PDE evaluations due to repeated failures, mainly caused by implementation issues or limitations in handling high-dimensional inputs. In particular, the pretrained transformer model for the End2end method was trained on synthetic equations with up to 10 features (Kamienny et al. 2022), which prohibits its usage on high-dimensional systems or systems with more than two state variables.
@@ 4.3 Limitations
Fourth, there is no established metric for quantifying equation fidelity. Standard metrics such as NMSE and symbolic complexity do not always reflect the symbolic correctness of discovered equations. For example, while GP methods achieved a relatively low NMSE on NS (Channel) dataset, the discovered equations do not capture the true dynamics. Without a robust metric that captures functional equivalence, e.g., using a discretized version of the Sobolev semi-norm for PDE systems (Schaback 2015), benchmarking results may obscure deeper algorithmic failures.
@@ Appendix A MDBench Pipeline
We release the MDBench pipeline as an open-source tool to support reproducible, extensible research in model discovery. The framework is modular by design — drawing inspiration from SRBench (La Cava et al. 2021) — and is structured to simplify the integration of new algorithms and datasets. All datasets are provided in a unified format to ensure cross-method compatibility and ease of adoption by the broader research community.
@@ Appendix B Datasets
Table A1 provides an overview of PDE systems in the benchmark. The newly generated datasets require the legacy FEniCs package (Alnæs et al. 2015), which can be readily installed on Ubuntu-based systems. Please refer to the README in the /scripts subfolder of the MDBench codebase for additional guidance on using FEniCS to generate the datasets, as well as the dataset-specific instructions below.
@@ B.1 Advection
The advection system, adopted from (Takamoto et al. 2022), is a simple linear system in 1D space with the equation
@@ B.2 Burgers
The Burgers equation, from (Takamoto et al. 2022), models the diffusion process in fluid dynamics as
@@
B.3 Korteweg–de Vries (KdV)
The KdV system, taken from (Brunton, Proctor, and Kutz 2016), models unidirectional propagation of small-amplitude waves in shallow water. The governing equation is given by
@@ B.3 Korteweg–de Vries (KdV)
B.4 Kuramoto Sivishinky (KS)
The KS equation is known for modeling the dynamics of various physical systems such as fluctuations in fluid films on inclines (Brunton, Proctor, and Kutz 2016). This equation includes a fourth-order diffusion term as a diffusive regularization of non-linear dynamics. The system is governed by the equation
@@ B.5 Nonlinear Schrödinger (NLS)
B.6 Reaction Diffusion
Diffusion reaction systems are used extensively to study pattern forming systems in physics (Brunton, Proctor, and Kutz 2016). One particular class of DR systems with nonlinear coupling terms are governed by the coupled PDE
@@ B.6 Reaction Diffusion
B.7 Advection Diffusion
The Advection Diffusion system, adopted from (Both et al. 2021), appears in heat transfer, pollution transport, and chemical physics, to name a few. We study the 2D Advection Diffusion system
@@ B.8 Heat (Solar)
The Heat (Solar) dataset from (Logg, Mardal, and Wells 2012) simulates the heating of the earth’s surface by solar radiation. For a given dimension $d\in\{1,2,3\}$, a rectangular domain $\omega$ is defined with $x_{d-1}=0$ representing the earth’s surface. Using default settings, the domain will be 1.5m in depth and 0.75m in width (for 2D and 3D problems). In the 3D case, we sample the first 20 time points.
@@ B.9 Heat (Laser)
The Heat (Laser) dataset from (Abali 2016) simulates the dynamics of heat transfer through a rigid body with a custom heat source. In this dataset, the heat supply comes from a moving laser beam that is used to heat a 3D plate in a concentrated manner. The power of the laser follows
@@
B.10 Navier–Stokes Channel (NS Channel)
The incompressible Navier–Stokes equations model the motion of viscous fluids. The NS Channel dataset, adopted from (Langtangen and Logg 2017), simulates a liquid traveling through a 2D channel along the $x$ direction. The dynamics follow the system of equations
@@ B.10 Navier–Stokes Channel (NS Channel)
B.11 Navier–Stokes Cylinder (NS Cylinder)
The NS Cylinder dataset is another example of a Navier-Stokes system as published in the FEATFLOW Benchmark Suite (Turek and Becker 1998). The system records the velocity and pressure generated by flow through a 2D rectangular channel 2.2 m in length and 0.41 m in width, with a 5 cm wide cylinder obstruction centered at $xy$-coordinates $(0.2,0.2)$. In this system, the fluid density $\rho=1$ and the kinematic velocity $\mu=0.001$. With a maximum velocity of 1.5 m/s, the system has a Reynolds number $Re=100$, which results in turbulent flow, with vortex shedding beyond the cylinder.
@@ B.11 Navier–Stokes Cylinder (NS Cylinder)
The unstructured mesh for the NS Cylinder dataset was generated using the mshr subpackage of FEniCS (Alnæs et al. 2015). The dataset simulates the system over 5 seconds, saving the data at 1,000 $Hz$ for a total of 5,000 timesteps. The dataset can be reproduced from the script fenics_navier_stokes_cylinder.py. Please note that the dataset must be converted into a structured mesh on a uniform grid prior to use for model discovery. The structured data can be generated from the unstructured version using the ns_cylinder_structured.py script. In our analysis, we subsample a rectangular grid in front of the cylinder (Figure A6) in space and $t\in[1,2],\Delta t=0.02$ in time.
@@
B.12 Diffusion-Reaction Cylinder
The Diffusion-Reaction Cylinder dataset utilizes the velocity component $\mathbf{v}$ of the solution to the NS Cylinder system to simulation an advection-diffusion-reaction system as detailed in (Langtangen and Logg 2017). Two chemical reagents $A$ and $B$ are carried by the fluid flow, i.e., advected, around the the cylinder and react to product chemical $C$, which is then disbursed downstream. The dynamical system of the chemical reactions is:
@@ B.12 Diffusion-Reaction Cylinder
Figure A7: Concentration of A in the Diffusion-Reaction Cylinder system at $t=199$.
Figure A8: Concentration of B in the Diffusion-Reaction Cylinder system at $t=199$.
Figure A9: Concentration of C in the Diffusion-Reaction Cylinder system at $t=199$.
The Diffusion-Reaction Cylinder is simulated for 5.0 seconds (matching the NS Cylinder velocity solution), saving the data at 100 $Hz$ for a total of 500 timesteps. The unstructured solution can be regenerated using the script fenics_diffusion_reaction_cylinder.py, from which the structured version can be generated using the diffusion_reaction_cylinder_structured.py script. We subsample the data both in time and space the same way as we did in the NS Cylinder system.
@@ Appendix C Methods and Hyperparameter Optimization
SINDy (Brunton, Proctor, and Kutz 2016) constructs a library of candidate functions (e.g., polynomials, trigonometric functions), evaluated on the state variables over time, and applies LASSO ($\ell_{1}$)-regularized regression to identify a sparse set of terms that best represent an ODE system. PDEFind (Rudy et al. 2017) extends SINDy to PDEs by evaluating the candidate library on a spatiotemporal grid.
WSINDy (Messenger and Bortz 2021) is another linear method for PDE discovery that avoids explicit pointwise derivative estimation by leveraging the weak form of the governing equations. ESINDy and EWSINDy (Fasel et al. 2022) extend SINDy and WSINDy, respectively, using bootstrap aggregation for improved robustness on ODE and PDE systems. All SINDy-based methods are implemented using the PySINDy library https://pysindy.readthedocs.io/en/latest/.
DeepMoD (Both et al. 2021) is a linear approach that integrates a neural network to simultaneously learn the underlying function, its derivatives, and the coefficients of a predefined candidate library. Bayesian (More et al. 2023) builds on PDEFind by incorporating variational Bayes and a spike-and-slab prior for sparse model selection. Its implementation is available at https://github.com/TapasTripura/Bayesian-Discovery-of-PDEs.
PySR (Cranmer 2023) is a genetic programming method with an optimized Julia backend, enabling high-performance model discovery through a parallel population-based evolve-simplify-optimize loop. See the official repository at https://github.com/MilesCranmer/PySR. Operon (Burlacu, Kronberger, and Kommenda 2020) is another efficient genetic programming method, implemented in C++ with vectorized operations and a compact linear tree representation. Its codebase is available at https://github.com/heal-research/operon.
EQL (Martius and Lampert 2016) is a deep learning-based method that embeds symbolic operations (e.g., addition, cosine) into the activation functions of a shallow neural network. After training, the network weights correspond to the coefficients of the discovered equation. We use the implementation available at https://github.com/cavalab/eql. uDSR (Landajuela et al. 2022) provides a unified framework that integrates mutiple approaches, from linear models to genetic programming, for model discovery. Its code is available at https://github.com/dso-org/deep-symbolic-optimization.
End2end (Kamienny et al. 2022) is a general-purpose model discovery approach based on a sequence-to-sequence Transformer architecture, pre-trained on large datasets of synthetic equations. The implementation can be found at https://github.com/facebookresearch/symbolicregression. ODEFormer (d’Ascoli et al. 2024) extends End2end to ODE systems. Unlike most generic model discovery methods, ODEFormer does not rely on derivative estimates and instead takes state trajectories and timestamps as input. Its implementation is available at https://github.com/sdascoli/odeformer.
@@ D.1 ODE Systems
Figure 10(a) presents aggregated performance of model discovery methods on all ODE systems, as measured by NMSE of predicted derivatives on the test set. While LM tend to discover less complex equations as the noise level increases, GP methods and ODEFormer exhibit the opposite trend. Figures A12 and A13 provide more detailed metrics on how different noise levels affect the quality of equations for each dataset. The dataset names are described in Table A4. For more details about the equations and initial values, refer to (d’Ascoli et al. 2024) or scripts/strogatz_ode.py in the MDBench project repository.
@@
LM and GP methods have a higher successful discovery rate compared to other evaluated methods. It can be observed that EQL is the most unstable method with the highest rate of failure. Both EQL and End2end were not able to accurately discover any of the ODE systems in the clean settings. The inaccuracy of End2end method likely stems from the lack of dynamical system examples in the datasets that the underlying model was pre-trained on. On the other hand, ODEFormer, while having a similar architecture to End2end, achieves a higher successful rediscovery rate. This suggests the necessity of including representative systems in the pre-training datasets for LSPT methods. A promising direction for improving the LSPT methods is to extend End2end and ODEFormer architectures to PDE systems, where a single transformer model is jointly trained on synthetic ODE and PDE systems.

### v2

@@ 4.2 Performance on PDE Data
Compared to ODEs, PDE systems present a more challenging testbed due to the presence of spatial derivatives and higher-order terms, which tend to amplify noise. Figure 2(b) shows the aggregate NMSE and complexity of representative methods over a selected subset of PDE datasets for which all methods completed successfully (refer to Figure 10(b) for all methods). The selected datasets include Advection, Burgers, KdV, KS, Heat (Solar) 1D, and AD. The primary reasons for a method’s failure to discover an underlying model include: a lack of support for higher-dimensional data in the published software, failing to finish training before the timeout threshold was reached, and runtime errors within the published codebases. EQL and End2end are excluded from the PDE evaluations due to repeated failures, mainly caused by implementation issues or limitations in handling high-dimensional inputs. In particular, the pretrained transformer model for the End2end method has been trained on synthetic equations with up to 10 features (kamienny2022end), which prohibits its usage on high-dimensional systems or systems with more than two state variables.
@@ 4.3 Limitations
Fourth, there is no established metric for quantifying equation fidelity. Standard metrics such as NMSE and symbolic complexity do not always reflect the symbolic correctness of discovered equations. For example, while GP methods achieved a relatively low NMSE on NS (Channel) dataset, the discovered equations do not capture the true dynamics. Without a robust metric that captures functional equivalence, e.g., using a discretized version of the Sobolev semi-norm for PDE systems (schaback2015computational), benchmarking results may obscure deeper algorithmic failures.
@@ Appendix A MDBench Pipeline
We release the MDBench pipeline as an open-source tool to support reproducible, extensible research in model discovery. The framework is modular by design — drawing inspiration from SRBench (la2021contemporary) — and is structured to simplify the integration of new algorithms and datasets. All datasets are provided in a unified format to ensure cross-method compatibility and ease of adoption by the broader research community.
@@ Appendix B Datasets
Table A1 provides an overview of PDE systems in the benchmark. The newly generated datasets require the legacy FEniCs package (fenics), which can be readily installed on Ubuntu-based systems. Please refer to the README in the /scripts subfolder of the MDBench codebase for additional guidance on using FEniCS to generate the datasets, as well as the dataset-specific instructions below.
@@ B.1 Advection
The advection system, adopted from (takamoto2022pdebench), is a simple linear system in 1D space with the equation
@@ B.2 Burgers
The Burgers equation, from (takamoto2022pdebench), models the diffusion process in fluid dynamics as
@@
B.3 Korteweg-de Vries (KdV)
The KdV system, taken from (brunton2016discovering), models unidirectional propagation of small-amplitude waves in shallow water. The governing equation is given by
@@ B.3 Korteweg-de Vries (KdV)
B.4 Kuramoto-Sivishinky (KS)
The KS equation is known for modeling the dynamics of various physical systems such as fluctuations in fluid films on inclines (brunton2016discovering). This equation includes a fourth-order diffusion term as a diffusive regularization of non-linear dynamics. The system is governed by the equation
@@ B.5 Nonlinear Schrödinger (NLS)
B.6 Reaction-Diffusion
Reaction diffusion systems are used extensively to study pattern forming systems in physics (brunton2016discovering). One particular class of DR systems with nonlinear coupling terms are governed by the coupled PDE
@@ B.6 Reaction-Diffusion
B.7 Advection-Diffusion
The Advection-Diffusion system, adopted from (both2021deepmod), appears in heat transfer, pollution transport, and chemical physics, to name a few. We study the 2D Advection-Diffusion system
@@ B.8 Heat (Solar)
The Heat (Solar) dataset from (logg2012automated) simulates the heating of the earth’s surface by solar radiation. For a given dimension $d\in\{1,2,3\}$, a rectangular domain $\omega$ is defined with $x_{d-1}=0$ representing the earth’s surface. Using default settings, the domain will be 1.5m in depth and 0.75m in width (for 2D and 3D problems). In the 3D case, we sample the first 20 time points.
@@ B.9 Heat (Laser)
The Heat (Laser) dataset from (abali2016computational) simulates the dynamics of heat transfer through a rigid body with a custom heat source. In this dataset, the heat supply comes from a moving laser beam that is used to heat a 3D plate in a concentrated manner. The power of the laser follows
@@
B.10 Navier-Stokes Channel (NS Channel)
The incompressible Navier-Stokes equations model the motion of viscous fluids. The NS Channel dataset, adopted from (langtangen2017solving), simulates a liquid traveling through a 2D channel along the $x$ direction. The dynamics follow the system of equations
@@ B.10 Navier-Stokes Channel (NS Channel)
B.11 Navier-Stokes Cylinder (NS Cylinder)
The NS Cylinder dataset is another example of a Navier-Stokes system as published in the FEATFLOW Benchmark Suite (featflow). The system records the velocity and pressure generated by flow through a 2D rectangular channel 2.2 m in length and 0.41 m in width, with a 5 cm wide cylinder obstruction centered at $xy$-coordinates $(0.2,0.2)$. In this system, the fluid density $\rho=1$ and the kinematic velocity $\mu=0.001$. With a maximum velocity of 1.5 m/s, the system has a Reynolds number $Re=100$, which results in turbulent flow, with vortex shedding beyond the cylinder.
@@ B.11 Navier-Stokes Cylinder (NS Cylinder)
The unstructured mesh for the NS Cylinder dataset was generated using the mshr subpackage of FEniCS (fenics). The dataset simulates the system over 5 seconds, saving the data at 1,000 $Hz$ for a total of 5,000 timesteps. The dataset can be reproduced from the script fenics_navier_stokes_cylinder.py. Please note that the dataset must be converted into a structured mesh on a uniform grid prior to use for model discovery. The structured data can be generated from the unstructured version using the ns_cylinder_structured.py script. In our analysis, we subsample a rectangular grid in front of the cylinder (Figure A6) in space and $t\in[1,2],\Delta t=0.02$ in time.
@@
B.12 Reaction-Diffusion (Cylinder)
The Reaction-Diffusion Cylinder dataset utilizes the velocity component $\mathbf{v}$ of the solution to the NS Cylinder system to simulation an advection-diffusion-reaction system as detailed in (langtangen2017solving). Two chemical reagents $A$ and $B$ are carried by the fluid flow, i.e., advected, around the the cylinder and react to product chemical $C$, which is then disbursed downstream. The dynamical system of the chemical reactions is:
@@ B.12 Reaction-Diffusion (Cylinder)
Figure A7: Concentration of A in the Reaction-Diffusion Cylinder system at $t=199$.
Figure A8: Concentration of B in the Reaction-Diffusion Cylinder system at $t=199$.
Figure A9: Concentration of C in the Reaction-Diffusion Cylinder system at $t=199$.
The Reaction-Diffusion Cylinder is simulated for 5.0 seconds (matching the NS Cylinder velocity solution), saving the data at 100 $Hz$ for a total of 500 timesteps. The unstructured solution can be regenerated using the script fenics_diffusion_reaction_cylinder.py, from which the structured version can be generated using the diffusion_reaction_cylinder_structured.py script. We subsample the data both in time and space the same way as we did in the NS Cylinder system.
@@ Appendix C Methods and Hyperparameter Optimization
SINDy (brunton2016discovering) constructs a library of candidate functions (e.g., polynomials, trigonometric functions), evaluated on the state variables over time, and applies LASSO ($\ell_{1}$)-regularized regression to identify a sparse set of terms that best represent an ODE system. PDEFind (rudy2017data) extends SINDy to PDEs by evaluating the candidate library on a spatiotemporal grid.
WSINDy (messenger2021weak) is another linear method for PDE discovery that avoids explicit pointwise derivative estimation by leveraging the weak form of the governing equations. ESINDy and EWSINDy (fasel2022ensemble) extend SINDy and WSINDy, respectively, using bootstrap aggregation for improved robustness on ODE and PDE systems. All SINDy-based methods are implemented using the PySINDy library https://pysindy.readthedocs.io/en/latest/.
DeepMoD (both2021deepmod) is a linear approach that integrates a neural network to simultaneously learn the underlying function, its derivatives, and the coefficients of a predefined candidate library. Bayesian (more2023bayesian) builds on PDEFind by incorporating variational Bayes and a spike-and-slab prior for sparse model selection. Its implementation is available at https://github.com/TapasTripura/Bayesian-Discovery-of-PDEs.
PySR (cranmer2023interpretable) is a genetic programming method with an optimized Julia backend, enabling high-performance model discovery through a parallel population-based evolve-simplify-optimize loop. See the official repository at https://github.com/MilesCranmer/PySR. Operon (burlacu2020operon) is another efficient genetic programming method, implemented in C++ with vectorized operations and a compact linear tree representation. Its codebase is available at https://github.com/heal-research/operon.
EQL (martius2016extrapolation) is a deep learning-based method that embeds symbolic operations (e.g., addition, cosine) into the activation functions of a shallow neural network. After training, the network weights correspond to the coefficients of the discovered equation. We use the implementation available at https://github.com/cavalab/eql. uDSR (landajuela2022unified) provides a unified framework that integrates multiple approaches, from linear models to genetic programming, for model discovery. Its code is available at https://github.com/dso-org/deep-symbolic-optimization.
End2end (kamienny2022end) is a general-purpose model discovery approach based on a sequence-to-sequence Transformer architecture, pretrained on large datasets of synthetic equations. The implementation can be found at https://github.com/facebookresearch/symbolicregression. ODEFormer (d2024odeformer) extends End2end to ODE systems. Unlike most generic model discovery methods, ODEFormer does not rely on derivative estimates and instead takes state trajectories and timestamps as input. Its implementation is available at https://github.com/sdascoli/odeformer.
@@ D.1 ODE Systems
Figure 10(a) presents aggregated performance of model discovery methods on all ODE systems, as measured by NMSE of predicted derivatives on the test set. While LM tend to discover less complex equations as the noise level increases, GP methods and ODEFormer exhibit the opposite trend. Figures A12 and A13 provide more detailed metrics on how different noise levels affect the quality of equations for each dataset. The dataset names are described in Table A4. For more details about the equations and initial values, refer to (d2024odeformer) or scripts/strogatz_ode.py in the MDBench project repository.
@@
LM and GP methods have a higher successful discovery rate compared to other evaluated methods. It can be observed that EQL is the most unstable method with the highest rate of failure. Both EQL and End2end were not able to accurately discover any of the ODE systems in the clean settings. The inaccuracy of End2end method likely stems from the lack of dynamical system examples in the datasets that the underlying model was pretrained on. On the other hand, ODEFormer, while having a similar architecture to End2end, achieves a higher successful rediscovery rate. This suggests the necessity of including representative systems in the pretraining datasets for LSPT methods. A promising direction for improving the LSPT methods is to extend End2end and ODEFormer architectures to PDE systems, where a single transformer model is jointly trained on synthetic ODE and PDE systems.